OpenStudy (thedj4jc):

What is the range of the function y = x2? all real numbers x ≥ 0 y ≥ 0

OpenStudy (freckles):

Do you know how to graph f(x)=x^2?

OpenStudy (freckles):

for range think y values

OpenStudy (thedj4jc):

No and it's not x to the second power, it's x times 2.

OpenStudy (freckles):

so its f(x)=2x?

OpenStudy (freckles):

ok well do you know how to graph f(x)=2x?

OpenStudy (thedj4jc):

No

OpenStudy (freckles):

Do you know that f(x)=mx+b is a linear function?

OpenStudy (freckles):

f(x)=mx+b where m is the slope and b is the y-intercept

OpenStudy (freckles):

you have f(x)=2x+0 so your y-intercept is? and your slope if ?

OpenStudy (freckles):

is*

OpenStudy (thedj4jc):

y is 0 and 2 is m???

OpenStudy (freckles):

yes

OpenStudy (freckles):

so use that info to graph your line

OpenStudy (freckles):

but you know your lines goes on forever and ever in either direction

OpenStudy (freckles):

So you actually already know the range but if you aren't convinced then I say still graph it

OpenStudy (thedj4jc):

Is it C, y >= 0?

OpenStudy (freckles):

Well it goes on in either direction forever and ever that just says the line starts above the x-axis and goes in the positive direction forever and ever

OpenStudy (freckles):

 this line goes on in either direction forever and ever